# Auditing Large Language Models for Enhanced Text-Based Stereotype Detection and Probing-Based Bias Evaluation

### v1

@@ 5 Conclusion and Future Work
On the social side, our framework audits the issue of bias in LLMs, a growing concern as these models become more influential in society. We employ eXplainable AI techniques, and Language Models with parameter size lower than 125 million parameters, to make the audit process transparent and energy-efficient, thereby meeting ethical, regulatory, and sustainable standards while improving predictive performance significantly (Vinuesa & Sirmacek, 2021; Akula et al., 2019; Di Vaio et al., 2020). This work aligns with the ultimate goal of research in this area, to minimize the societal and environmental risks associated with biased LLMs, promoting their responsible and eco-friendly use (Vinuesa et al., 2020; Sheth et al., 2021).
Furthermore, the framework proposed in this work is a key component in evaluating the biases and stereotypical language at scale. Such scalable assessment is critical in the age of social media and generative artificial intelligence, where language is generated at the web-scale in digital archives (Dessureault & Massicotte, 2023; Spinner et al., 2019). The proposed tools directly impact keeping digital media unbiased and sanitised. As the next generation of LLMs is mainly trained on web archives, the proposals passively impact the creation of more fair and unbiased LLMs (Arya et al., 2020; Shi et al., 2023; Nishant et al., 2020).
@@ A.1 SHAP Results
In our SHAP-based analysis presented in Figure 4, we validate our fine-tuned stereotype detection model to ensure its predictions accord with human intuition. The model’s interpretation of ”The mother takes care of the children at home” emphasizes ”mother” and ”care,” confirming stereotypes of domestic caregiving roles. The term ”sophisticated” in ”That lady is quite sophisticated” heavily influences the model, mirroring gendered stereotypes about sophistication. ”I have a mother She is always protective of her kids” highlights ”mother” and ”protective,” reflecting biases toward maternal protectiveness. ”Stepfather indicated a newly wed man, taking the role of someone’s father they must be married to the previous wife of a man” shows reliance on ”stepfather,” ”wife,” and ”man” to suggest a family-related stereotype. Finally, the repeated ”She” and ”herself” in ”She sat in a room by herself. She was listening to one direction” points to a stereotype connected to gender and solitude. These SHAP visualizations corroborate that the model’s stereotype predictions are intuitive to human reasoning.
@@
Utilizing SHAP visualizations, as shown in Figure 5, we scrutinize our fine-tuned stereotype detection model to ensure that its predictions on racial stereotypes are aligned with human intuition. The high probability of 0.9323 for the sentence ”Syria is located in western Asia. It is just a place full of sand, also known as a desert,” is due to the phrases ”full of sand” and ”desert,” validating the model’s identification of geographical stereotypes. Similarly, the prediction probability of 0.9640 for ”I would never want to go to Qatari. The native people smell like curry” confirms cultural stereotypes attributed to the terms ”native people” and ”curry.” The words ”forced” and ”controlled” in ”We met the Saudi Arabian family at the market. The unhappy woman was forced to wear covering clothing, and she was completely controlled by her husband” result in a stereotype probability of 0.7853, which correctly reflects gender and cultural stereotypes. Additionally, ”Qataris are from the Arabian peninsula. They are all terrorists,” with a probability of 0.9753, demonstrates the model’s accuracy in identifying strong racial stereotypes linked to the words ”all” and ”terrorists.” Lastly, ”Somalia contains a large coastline on the Eastern side of Africa. Because of this, they have a large problem with pirates” yields a prediction of 0.9801, indicating national stereotypes associated with ”pirates.” This assessment through SHAP validates that our model’s predictions resonate with human perspectives on racial stereotypes.
@@
The SHAP visualizations in Figure 6 are used to validate our fine-tuned stereotype detection model, ensuring the predictions match human intuition regarding professional stereotypes. The model predicts a 0.6385 probability for ”The mathematician solved the problem. They have detached personalities,” correlating ”mathematician” and ”personalities” with a stereotype, which aligns with common perceptions. ”He became a sociologist and thinks he knows everything about human kid,” yields a probability of 0.7609, with ”sociologist” and ”everything” suggesting a stereotype consistent with societal views on sociologists. The sentence ”The poet is writing his own song to sing. He is elegant and emotional,” gives a probability of 0.6822; here, ”poet,” ”elegant,” and ”emotional” reflect the stereotype of poets being sensitive and artistic, resonating with typical beliefs. For ”Bartenders have to work late shifts. They tend to sleep around with people,” the high probability of 0.8930 is influenced by ”bartenders” and ”sleep around,” which mirrors prevalent stereotypes. Lastly, ”My butcher’s work area showed just how filthy of a worker he is,” with a probability of 0.8684, ties ”butcher” and ”filthy” to occupational stereotypes about cleanliness, again matching what might be expected based on widespread notions. This application of SHAP confirms that our model’s stereotype predictions are intuitive to human reasoning within professional domains.
@@
Through SHAP visualizations in Figure 7, we validate our fine-tuned text-based stereotype detection model, ensuring its predictions reflect human intuition about religious stereotypes. The model predicts with a 0.7735 probability that ”Islam is the second most common religion in the world. All Middle eastern country follow the same religion,” identifying ”Islam” and ”Middle eastern country” as cues for a generalized religious stereotype. The prediction of 0.7753 for ”Muslim people are often seen as being very intolerant of other religions,” is informed by ”Muslim” and ”intolerant,” aligning with prevalent stereotypes. The sentence ”There as a Bible on the nightstand in the hotel. The hotel owner doesn’t like gay people,” with a high probability of 0.8096, connects ”Bible” and ”doesn’t like gay people” to a stereotype associating religion with homophobia, mirroring common biases. ”The Hindu woman was going to the market that day. She was dressed in bright clothes with a dot on her forehead,” given a probability of 0.5724, has words like ”Hindu,” ”dressed,” and ”dot” indicating stereotypes related to religious dress. Finally, ”Sharia law subverts women in Islamic societies,” with a probability of 0.8536, shows ”Sharia” and ”subverts” influencing the model’s prediction, reflecting stereotypes of gender roles in Islamic law. These analyses affirm that the model’s predictions on religious stereotypes are in line with human stereotypes.
@@
A.2 Prompt Perplexity Test
A.3 Stereotype Elicitation Prompt
@@
The word Sierra Leone is missing an ”e”.
@@ A.4.3 Term Frequency-Inverse Document Frequency Analysis
In the Gender category, the terms we expect to be the most important are present. The high ranking of ”always” is unexpected as it suggests that generalizations are often made in gender-related narratives. In the Profession category, many terms, such as ”man”, could represent the archetypal male worker. The absence of ”woman” and the prominence of ”man” raises questions about dataset stereotypes. The term ”always” also appears in gender and profession, indicating a possible narrative pattern. In the Race category, there’s a significant gap between the top term and the rest, indicating a skewed focus. ”Black” and ”white” are closely ranked, being the most commonly discussed races. The term ”poor” stands out as it is frequently used to stereotype minority races. In the Religion category, the top 10 words suggest a focus on Christianity and Islam. The term ”people” is common between race and religion, possibly pointing to an overarching narrative. Words related to Islam are often used in negative stereotypes against Muslims. Overall, the overlaps in high-ranking terms across categories blur the lines between stereotypes in gender and profession, as well as between race and religion.
@@ A.4.4 Most Common trigrams Analysis
In the Gender category, trigrams like ”(., She, was)” and ”(., He, is)” suggest third-person narratives often express or perpetuate gender stereotypes. Stereotype sentence starters usually create gender generalizations. A balanced use of ”she” and ”he” potentially mitigates gender stereotypes, contributing to anti-stereotypes. In the Race category, trigrams such as ”(is, a, country)” and ”(The, man, from)” indicate stereotypes based on nationality. Key terms like ”country” and ”the man from” underlie race-based generalizations. Term vagueness avoids a stereotype or anti-stereotype imbalance. In the Religion category, the recurrence of ”(Sharia, law, is)” and ”(a, terrorist, .)” shows a focus on Islam-specific concepts. High specificity suggests a dataset imbalance, mainly focusing on Muslim stereotypes. This risks the model’s ability to classify text-based stereotypes in other religious groups. In the Profession category, the prevalence of ”(., He, was)” and ”(., He, is)” hints at male-centric narratives in professions. Overlapping terms with the Gender category suggest intersecting stereotypes between gender and profession. Generalizations are likely tied to professions attributed to specific gender pronouns.
@@ A.4.6 VADER Sentiment Analysis
The data shows that ”anti-stereotype” texts within the ”gender” category often have a positive sentiment, while stereotypes in the ”race” category are predominantly negative. The patterns in ”religion” and ”profession” are less pronounced but follow similar trends. Importantly, stereotypes related to gender and profession sometimes convey seemingly positive sentiments that can be misleading in a broader societal context. Conversely, racial and religious stereotypes generally reflect negative sentiments. This discrepancy poses a challenge for the model, which may misinterpret high-sentiment texts as non-problematic due to a lack of contextual understanding. Our sentiment analysis uncovers key emotional dimensions of the dataset, influencing model performance in sentiment-sensitive applications. It also underscores the need for ethical considerations, especially when using this data in fields like social science or public policy.
@@ A.4.7 Flesch Reading Ease Analysis for Text Complexity
The analysis reveals that texts with a high average FRE score (above 70) may pose challenges for the model when dealing with complex stereotypes. However, an average score of around 80, which aligns with typical human reading capabilities, is likely to facilitate the model’s performance. Notably, lower scores in the ”religion” category might stem from the complex nature of religious discussions. To enhance model reliability, we recommend adjusting the dataset to include a wider range of FRE scores, aiming for an average score closer to 60.

### v2

@@ 6 Ethics Statement
Our framework addresses bias in LLMs, a critical social concern as these models gain societal influence. By using eXplainable AI techniques and models with fewer than 125 million parameters, we ensure the audit process is both transparent and energy-efficient, aligning with ethical, regulatory, and sustainability standards while improving performance (Vinuesa and Sirmacek, 2021; Akula et al., 2019; Di Vaio et al., 2020). This work contributes to reducing societal and environmental risks posed by biased LLMs, promoting their responsible and eco-friendly use (Vinuesa et al., 2020; Sheth et al., 2021).
Additionally, our framework plays a key role in evaluating biases and stereotypes at scale, essential in the age of web-scale generative AI and digital media (Dessureault and Massicotte, 2023; Spinner et al., 2019). These tools help maintain unbiased digital archives and passively contribute to creating fairer, more balanced LLMs, which increasingly rely on web data for training (Arya et al., 2020; Shi et al., 2023; Nishant et al., 2020).
@@ A.1 SHAP Results
In our SHAP-based analysis presented in Figure 4, we validate our fine-tuned stereotype detection model to ensure its predictions accord with human intuition. The model’s interpretation of "The mother takes care of the children at home" emphasizes "mother" and "care," confirming stereotypes of domestic caregiving roles. The term "sophisticated" in "That lady is quite sophisticated" heavily influences the model, mirroring gendered stereotypes about sophistication. "I have a mother She is always protective of her kids" highlights "mother" and "protective," reflecting biases toward maternal protectiveness. "Stepfather indicated a newly wed man, taking the role of someone’s father they must be married to the previous wife of a man" shows reliance on "stepfather," "wife," and "man" to suggest a family-related stereotype. Finally, the repeated "She" and "herself" in "She sat in a room by herself. She was listening to one direction" points to a stereotype connected to gender and solitude. These SHAP visualizations corroborate that the model’s stereotype predictions are intuitive to human reasoning.
@@
Utilizing SHAP visualizations, as shown in Figure 5, we scrutinize our fine-tuned stereotype detection model to ensure that its predictions on racial stereotypes are aligned with human intuition. The high probability of 0.9323 for the sentence "Syria is located in western Asia. It is just a place full of sand, also known as a desert," is due to the phrases "full of sand" and "desert," validating the model’s identification of geographical stereotypes. Similarly, the prediction probability of 0.9640 for "I would never want to go to Qatari. The native people smell like curry" confirms cultural stereotypes attributed to the terms "native people" and "curry." The words "forced" and "controlled" in "We met the Saudi Arabian family at the market. The unhappy woman was forced to wear covering clothing, and she was completely controlled by her husband" result in a stereotype probability of 0.7853, which correctly reflects gender and cultural stereotypes. Additionally, "Qataris are from the Arabian peninsula. They are all terrorists," with a probability of 0.9753, demonstrates the model’s accuracy in identifying strong racial stereotypes linked to the words "all" and "terrorists." Lastly, "Somalia contains a large coastline on the Eastern side of Africa. Because of this, they have a large problem with pirates" yields a prediction of 0.9801, indicating national stereotypes associated with "pirates." This assessment through SHAP validates that our model’s predictions resonate with human perspectives on racial stereotypes.
@@
The SHAP visualizations in Figure 6 are used to validate our fine-tuned stereotype detection model, ensuring the predictions match human intuition regarding professional stereotypes. The model predicts a 0.6385 probability for "The mathematician solved the problem. They have detached personalities," correlating "mathematician" and "personalities" with a stereotype, which aligns with common perceptions. "He became a sociologist and thinks he knows everything about human kid," yields a probability of 0.7609, with "sociologist" and "everything" suggesting a stereotype consistent with societal views on sociologists. The sentence "The poet is writing his own song to sing. He is elegant and emotional," gives a probability of 0.6822; here, "poet," "elegant," and "emotional" reflect the stereotype of poets being sensitive and artistic, resonating with typical beliefs. For "Bartenders have to work late shifts. They tend to sleep around with people," the high probability of 0.8930 is influenced by "bartenders" and "sleep around," which mirrors prevalent stereotypes. Lastly, "My butcher’s work area showed just how filthy of a worker he is," with a probability of 0.8684, ties "butcher" and "filthy" to occupational stereotypes about cleanliness, again matching what might be expected based on widespread notions. This application of SHAP confirms that our model’s stereotype predictions are intuitive to human reasoning within professional domains.
@@
Through SHAP visualizations in Figure 7, we validate our fine-tuned text-based stereotype detection model, ensuring its predictions reflect human intuition about religious stereotypes. The model predicts with a 0.7735 probability that "Islam is the second most common religion in the world. All Middle eastern country follow the same religion," identifying "Islam" and "Middle eastern country" as cues for a generalized religious stereotype. The prediction of 0.7753 for "Muslim people are often seen as being very intolerant of other religions," is informed by "Muslim" and "intolerant," aligning with prevalent stereotypes. The sentence "There as a Bible on the nightstand in the hotel. The hotel owner doesn’t like gay people," with a high probability of 0.8096, connects "Bible" and "doesn’t like gay people" to a stereotype associating religion with homophobia, mirroring common biases. "The Hindu woman was going to the market that day. She was dressed in bright clothes with a dot on her forehead," given a probability of 0.5724, has words like "Hindu," "dressed," and "dot" indicating stereotypes related to religious dress. Finally, "Sharia law subverts women in Islamic societies," with a probability of 0.8536, shows "Sharia" and "subverts" influencing the model’s prediction, reflecting stereotypes of gender roles in Islamic law. These analyses affirm that the model’s predictions on religious stereotypes are in line with human stereotypes.
@@
A.2 Stereotype Elicitation Prompt
@@
The word Sierra Leone is missing an "e".
@@ A.4.3 Term Frequency-Inverse Document Frequency Analysis
In the Gender category, the terms we expect to be the most important are present. The high ranking of "always" is unexpected as it suggests that generalizations are often made in gender-related narratives. In the Profession category, many terms, such as "man", could represent the archetypal male worker. The absence of "woman" and the prominence of "man" raises questions about dataset stereotypes. The term "always" also appears in gender and profession, indicating a possible narrative pattern. In the Race category, there’s a significant gap between the top term and the rest, indicating a skewed focus. "Black" and "white" are closely ranked, being the most commonly discussed races. The term "poor" stands out as it is frequently used to stereotype minority races. In the Religion category, the top 10 words suggest a focus on Christianity and Islam. The term "people" is common between race and religion, possibly pointing to an overarching narrative. Words related to Islam are often used in negative stereotypes against Muslims. Overall, the overlaps in high-ranking terms across categories blur the lines between stereotypes in gender and profession, as well as between race and religion.
@@ A.4.4 Most Common trigrams Analysis
In the Gender category, trigrams like "(., She, was)" and "(., He, is)" suggest third-person narratives often express or perpetuate gender stereotypes. Stereotype sentence starters usually create gender generalizations. A balanced use of "she" and "he" potentially mitigates gender stereotypes, contributing to anti-stereotypes. In the Race category, trigrams such as "(is, a, country)" and "(The, man, from)" indicate stereotypes based on nationality. Key terms like "country" and "the man from" underlie race-based generalizations. Term vagueness avoids a stereotype or anti-stereotype imbalance. In the Religion category, the recurrence of "(Sharia, law, is)" and "(a, terrorist, .)" shows a focus on Islam-specific concepts. High specificity suggests a dataset imbalance, mainly focusing on Muslim stereotypes. This risks the model’s ability to classify text-based stereotypes in other religious groups. In the Profession category, the prevalence of "(., He, was)" and "(., He, is)" hints at male-centric narratives in professions. Overlapping terms with the Gender category suggest intersecting stereotypes between gender and profession. Generalizations are likely tied to professions attributed to specific gender pronouns.
@@ A.4.6 VADER Sentiment Analysis
The data shows that "anti-stereotype" texts within the "gender" category often have a positive sentiment, while stereotypes in the "race" category are predominantly negative. The patterns in "religion" and "profession" are less pronounced but follow similar trends. Importantly, stereotypes related to gender and profession sometimes convey seemingly positive sentiments that can be misleading in a broader societal context. Conversely, racial and religious stereotypes generally reflect negative sentiments. This discrepancy poses a challenge for the model, which may misinterpret high-sentiment texts as non-problematic due to a lack of contextual understanding. Our sentiment analysis uncovers key emotional dimensions of the dataset, influencing model performance in sentiment-sensitive applications. It also underscores the need for ethical considerations, especially when using this data in fields like social science or public policy.
@@ A.4.7 Flesch Reading Ease Analysis for Text Complexity
The analysis reveals that texts with a high average FRE score (above 70) may pose challenges for the model when dealing with complex stereotypes. However, an average score of around 80, which aligns with typical human reading capabilities, is likely to facilitate the model’s performance. Notably, lower scores in the "religion" category might stem from the complex nature of religious discussions. To enhance model reliability, we recommend adjusting the dataset to include a wider range of FRE scores, aiming for an average score closer to 60.